# DocFinQA: A Long-Context Financial Reasoning Dataset

### 12 Jan 2024 v1

@@ 4 Retrieval-based QA Evaluation
We evaluate these models using hit rate (HR@k). We compute the HR using target $c^{*}$ on the test set of DocFinQA. The results are shown in Figure 3. The Finetuned ColBERT model yields the highest HR, followed by the Original ColBERT model. The finetuned ColBERT model yields an average improvement of 91% HR over Sentence-BERT. We obtain a maximum 0.35 (HR@1) and 0.55 (HR@3).
Question Answering Task We formulate the QA task as a few-shot in-context learning task Brown et al. (2020). Due to the limited context length, the context of each demonstration contains a single golden chunk, while the context of the query question may contain up to $k$ chunks. A sample prompt template is given in Figure 10.
Table 2 reports the performance of 11 state-of-the art models with low context length. Larger models perform better than smaller models, as expected (e.g., MPT 30B vs MPT 7B). Models trained on code yield higher accuracy than non-code models (e.g., CodeLlama vs Llama). Models with additional supervised finetuning (e.g., LLama 2/7B+SFT) and instruction tuning (e.g., GPT-3.5) are the best among the examined models. Notably, Mistral 7B outperforms several larger models, although it lags behind Llama 2/7B+SFT and GPT-3.5.
Finetuned ColBERT is the best retrieval model in all but one setting. It yields a marginal but consistent improvement over the Original ColBERT, and a large improvement over Sentence BERT and OpenAI ADA.
Model/Size ColB. SentB. ADA FT ColB.
Falcon/7B   2.31   0.30   1.21   1.82
MPT/7B   4.63   2.71   3.83   4.77
MPT/30B 17.32 11.11 12.12 18.07
Llama 2/7B 13.49 8.86 11.11 13.50
Llama 2/13B 18.73 12.68 14.94 19.13
CodeLlama/7B 15.60 12.18 15.15 16.75
CodeLlama/13B 19.10 13.79 18.78 21.01
Mistral/7B 23.16 14.90 21.51 24.97
Llama 2/7B+SFT 32.93 24.77 34.34 36.14
GPT-3.5/- 41.59 33.78 36.43 42.64
Table 2: Performance of the LLMs on DocFinQA test set. For each row, the best performance among all retrieval models is in bold. The fewshot setting is selected based on the best performance on the development set (See Appendix I).
5 Retrieval-free QA Evaluation
Recently developed LLMs can handle context lengths over 100k tokens. We explore the promise of these models for retrieval-free long document QA using the DocFinQA dataset.
Since the average size of documents in our data is still longer than the context window, we adopt different strategies for answering questions over short (fewer than 100K tokens) and longer documents. For short documents, we provide examples from FinQA dataset (i.e., with  700 words of context) as a few-shot prompt followed by the whole target document and question. For longer documents, we adapt the S2A strategy proposed in Weston and Sukhbaatar (2023), which summarizes relevant information from the document before answering the question. For each 100K chunk of the long document, we include question/context pairs from FinQA as examples of information to extract for a summary. We then input the generated summary to the model with the same prompt used with short documents. Due to budget constraints, we sub-sample 400 questions from the test set for conducting experiments with GPT4. We refer to this set as the “test subsample”. In our test set, we maintain an equal number of short and longer documents (i.e., 200 datapoints of each type).
Model w/ Retrieval Accuracy
GPT-4 No 34.8
    Short-only No 46.5
    Long-only w/ S2A No 23.0
    Long-only w/ Interactive No 20.0
Mistral/7B Yes 28.4
LLama 2/7B+SFT Yes 28.6
GPT-3.5 Yes 42.4
GPT-4 Yes 51.5
Table 3: Performance of LLMs on DocFinQA test set.
Table 3 shows a comparison of GPT-4 performance on the test subsample when given full document context. The performance on short documents is twice as good as over longer documents, but is still less than 50% accurate. This indicates that there is much room for improvement with document-level financial QA.
The difference between long and short document accuracy also implies that better methods are needed for combining information across multiple calls to a document-processing LM. In this work, we explored two options: the S2A setup described above, and an Iterative refinement approach where the output program is written and refined as the LLM processes each 100k section of the document. Here, a temporary answer program (initially “None”) is passed along with the input each time the LLM is called. This method proved slightly less performant than S2A, achieving 20% accuracy compared to 23% on the longer-document subset, respectively.
6 Conclusion
This paper introduces a highly practical setting for grounded document-level question answering over financial reports. Previous work provides select content, whereas we incorporate the full financial report — averaging 123k words — for each financial question. Our findings reveal that this more realistic setting presents a significantly more difficult challenge, thereby opening new avenues for research in quantitative financial question answering.
Appendix A Impact of Data Selection
@@
Appendix B Golden Chunk Position
@@
Appendix C Model Details
@@
We also included the Llama 2/7B + SFT that was finetuned on the training set of DocFinQA with golden chunk from FinQA ($c^{golden}$). The finetuning process takes 3 epochs with a batch size of 32. We use the context provided by the FinQA dataset as the input due to the limited maximum token length of the model. The maximum token length is set to 2048. The model is finetuned on 8 x Nvidia A100-80GB GPUs. We use AdamW optimizer with learning rate of 2e-6.
Appendix D Parsing SEC Filings
@@
Appendix E ColBERT Finetuning
@@
Appendix F Code Conversion
@@
Appendix G Few-shot Settings
@@
Appendix H Detailed performance of retrieval method
@@
Appendix I Detailed performance on dev set
@@
Appendix J Prompt examples
Table 5 and J shows examples of context, question, and model output in FinQA and DocFinQA.
@@
Table 5: Example from FinQA dataset. The context provided here has been formatted from on the original dataset values

### 29 Feb 2024 v2

@@ Appendix B Retrieval Performance
Appendix C Impact of Data Selection
@@
Appendix D Golden Chunk Position
@@
Appendix E Model Details
@@
We also included the Llama 2/7B + SFT that was finetuned on the training set of DocFinQA with golden chunk from FinQA ($c^{golden}$). The finetuning process takes 3 epochs with a batch size of 32. We use the context provided by the FinQA dataset as the input due to the limited maximum token length of the model. The maximum token length is set to 2048. The model is finetuned on 8 x Nvidia A100-80GB GPUs. We use AdamW optimizer with learning rate of 2e-6. The training process takes 4 hours to complete.
Appendix F SEC Filing Collection
Each data point in the FinQA dataset consists of a document identification field as shown in Table 5. This field is made up of 3 sections separated by a forward slash. The first is a string called company ticker symbol, the second refers to the year in which this document was filed and the third is the page number in the document where the answer can be found.
Downloading the right 10-K filing from SEC begins with identifying the company code from the company ticker symbol. For example, C/2017/page_328.pdf-1 in FinQA maps to the CITIGROUP INC with company code 831001. This mapping is obtained from the official file released by SEC which can be found here https://www.sec.gov/file/company-tickers. We automatically generate a URL using the company code obtained. From the SEC website, either filings are downloaded as TXT, HTML or XBRL using the generated URL. At this stage, approximately 6.5% (or 543) data points which is corresponds to approximately 9.4% (or 17) documents were dropped, either due to lack of mapping or nonavailability of older documents. Further, conversion of the downloaded files to PDF caused a loss of 117 data points (19 unique documents) due to formatting issues.
Appendix G Parsing SEC Filings
Since each filing contains many tables, maintaining the structure and order during extraction is critical for numerical reasoning. We convert each HTML-formatted filing to PDF format and use a finance-specific PDF extractor to parse the filing into markdown format. This process ensures that: (i) our dataset is grounded in the relevant financial documentation and (ii) all the tables in the filings are parsed with high precision into a consistent format without any HTML-tag noise.
@@ Appendix G Parsing SEC Filings
Appendix H ColBERT Finetuning
@@
Appendix I Code Conversion
@@
Appendix J Few-shot Settings
@@
Appendix K Detailed performance of retrieval method
@@
Appendix L Detailed performance on dev set
@@
Appendix M Prompt examples
Table 5 and M shows examples of context, question, and model output in FinQA and DocFinQA.
ID: C/2017/page_328.pdf-1
@@
Table 5: Example from FinQA dataset. The context provided here has been formatted from the original dataset values.